# Improving Requirements Classification with SMOTE-Tomek Preprocessing

### v1

@@ Abstract
Requirement Engineering, Class Imbalance, SMOTE-Tomek, Imbalanced Datasets, Requirement Classification, Natural Language Processing, Synthetic Oversampling, Text-Based Classification, Supervised Learning.
I Introduction
@@
Requirements engineering (RE) is a cornerstone of the software development lifecycle, translating stakeholder needs into actionable system specifications [1, 2]. Accurate and efficient requirement classification is essential to ensure project clarity, prioritization, and goal alignment. Requirements are typically categorized into functional, non-functional, and subtypes [3, 4, 5, 6]. A survey revealed that over 60% of failed projects neglected non-functional requirements, underscoring the critical importance of effective classification [7]. Despite its significance, requirement classification remains a largely manual process, prone to inconsistencies, inefficiencies, and scalability challenges [8].
@@
Over the years, researchers have explored various approaches to automate requirement classification. Early rule-based systems were among the first attempts, offering interpretable but inflexible solutions that were labor-intensive to maintain [9]. The advent of machine learning (ML) brought more adaptive models, demonstrating moderate success, particularly in structured datasets [10, 3, 6]. However, these methods often face challenges with class imbalance, a prevalent issue in real-world datasets, including the widely used PROMISE dataset for requirements engineering [11]. This imbalance skews model predictions toward majority classes, undermining performance on minority classes.
@@
Over the past decade, deep learning (DL) has achieved transformative breakthroughs across multiple domains, including natural language processing (NLP), computer vision, and time series analysis. These advancements are primarily driven by the unprecedented representational power of deep neural networks, which leverage millions, and in some cases billions, of trainable parameters to extract and model intricate, high-dimensional patterns from large-scale datasets [12].
@@
Equally significant are the advances in time series analysis, where DL models have demonstrated remarkable capabilities in tasks such as motion sensing classification and physical quantity estimation [13, 14, 15, 16]. These innovations highlight DL’s ability to capture temporal dependencies and complex properties in sequential data.
@@
In computer vision, architectures such as ResNet [17] and Vision Transformers (ViT) [18] have revolutionized image classification, object detection, and semantic segmentation, achieving state-of-the-art accuracy while enabling applications that were previously unattainable.
@@
In NLP, DL models such as BERT [19] and GPT-3 [20] have set new standards in tasks like language translation, text summarization, and question answering, surpassing traditional methods and enabling more nuanced understanding and generation of human language [21, 22, 23, 24, 25]. These models excel in NLP tasks due to their ability to capture contextual relationships within text using self-attention mechanisms. They are particularly effective at modeling semantic intricacies in textual data, enabling strong performance in complex classification tasks. However, these methods require not only substantial computational resources for fine-tuning and inference but also access to massive amounts of high-quality data to achieve their full potential. This reliance on extensive datasets and powerful hardware often makes them less practical for small-scale or resource-constrained projects, where data availability and computational capacity may be limited.
@@
To address these challenges, this study integrates the Synthetic Minority Oversampling Technique (SMOTE) [26] with Tomek Links (SMOTE-Tomek) [27], providing a robust preprocessing solution for imbalanced datasets. SMOTE oversamples minority classes by generating synthetic samples through interpolation, effectively enhancing the diversity of underrepresented classes while mitigating overfitting, a common issue with random oversampling methods. Complementing this, Tomek Links identifies and removes noisy or borderline samples, improving class separability by reducing overlapping data points between classes. This dual-action preprocessing strategy results in cleaner and more balanced training datasets, enabling ML models to better learn decision boundaries for minority classes.
@@
In RE classification, SMOTE-Tomek addresses both class imbalance and noise, making it particularly suitable for the PROMISE dataset, where minority classes are significantly underrepresented. While previous studies have demonstrated the efficacy of SMOTE-Tomek for addressing class imbalance in the PROMISE dataset [28], this research extends its application by integrating it into a stratified K-fold cross-validation framework. This ensures the validation folds remain untouched by resampling, preserving the original class distribution and enabling rigorous evaluation of the model’s generalization capabilities. Furthermore, this study explores several classical ML models, evaluates their performance on the PROMISE dataset, and highlights the potential of lightweight solutions for efficient requirement classification.
@@ II Learning Method
II-A Dataset
The study utilizes the expanded PROMISE dataset, a diverse collection of 969 software requirements extracted from Software Requirements Specification (SRS) documents using the Google search engine [11]. The dataset contains 444 functional requirements (45.8%) and 525 non-functional requirements (54.2%), distributed across 12 distinct categories. These categories represent a spectrum of requirements in software engineering, such as Security (SE), Usability (US), Portability (PO), and Performance (PE). Despite its comprehensiveness, the dataset exhibits notable class imbalance. For example, the Portability category is severely underrepresented, comprising only 12 requirements (1.24%), while the most prevalent non-functional class, Security, includes 125 requirements (12.9%). This disparity is further amplified in the broader classification between functional and non-functional requirements, which affects the performance and generalizability of ML models. Figure 1 illustrates the distribution of the dataset across all requirement types, highlighting the inherent imbalance.
@@ II-A Dataset
Table I provides examples of several requirements and their types from the PROMISE dataset.
TABLE I: Several Requirements and Their Types
From The PROMISE Dataset
@@
Users should be able to access their streaming movies in under 2 clicks after logging into the website. Usability (US)
@@
II-B Pre-Processing
Transforming unstructured text into a numerical representation suitable for ML algorithms is a fundamental preprocessing step in many NLP tasks. The PROMISE dataset comprises individual sentences, rather than full documents. The common Term Frequency-Inverse Document Frequency (TF-IDF) remains a highly effective vectorization technique [29, 30]. TF-IDF quantifies the importance of a term within a sentence while considering its prevalence across the entire dataset of sentences. By emphasizing terms that are frequent within a specific sentence but rare across the dataset, TF-IDF captures features that are contextually meaningful and relevant to the classification task. Despite the brevity of the textual units, TF-IDF preserves the semantic significance of terms.
II-C Class Imbalance Challenge
Class imbalance poses a critical challenge in training ML models, as they tend to prioritize majority classes, resulting in poor generalization and suboptimal performance on minority classes [31]. Such biases can significantly impact the usability and reliability of classification models in practical applications, where correctly identifying minority classes is often essential. Balancing the training set ensures that the model learns to give equal attention to all classes, preventing it from being overly biased toward majority classes.
II-D SMOTE-Tomek
@@
In practical applications of NLP tasks, where textual data is complex and often imbalanced, SMOTE-Tomek emerges as a robust and effective solution. The synthetic examples generated by SMOTE are not direct textual entities but instead constitute feature vectors that represent the underlying characteristics of the minority class, as presented in Table II. These vectors serve as abstract representations, enabling the model to internalize the structural properties of underrepresented classes without requiring additional annotated textual data.
@@
TABLE II: Examples of Requirements and Types Resampling Using SMOTE-Tomek
Requirement Type
based, established, lead, operate, physical, process, product, run, server, service, shall, structure, washing, web O
changes, edit, information, just, modify, normal, personal, read, securely, server, transmission, transmitted, users SE
entry, shown F
initiator, mediator, messages, request, response, shall, user F
accommodate, compensatory, data, failures, fault, items, large, product, recovery, robust, routing, shall, technique, tolerance, transaction, using FT
explanatory, intuitive, self, shall US
control, details, employees, site, view F
browsers, ce, compatible, environments, expected, interface, major, operating, order, palm, product, run, standards, systems, usable, variety, web, wide, windows PO
800x600, 95, appropriately, correct, corrected, display, feel, higher, implementation, incorrect, intranet, look, notification, pages, prior, provide, remaining, resolutions, shall, uniform, web, week LF
able, alongside, days, environment, function, installed, java, operating, product, runtime, server, shall, software O
II-E Stratified K-fold Cross-Validation Method
@@
II-F Trainig Algorithm
The proposed method, as outlined in Algorithm 1, emphasizes the integration of stratified K-fold cross-validation to ensure class proportions are preserved across folds. Importantly, SMOTE-Tomek is applied solely to the training folds to prevent data leakage, thereby preserving the integrity of validation metrics. The algorithm also includes final training on the entire balanced dataset to fully utilize the available data. Table III presents the notations and their corresponding definitions as utilized in the algorithm.
TABLE III: Notations and Definitions
@@ II-F Trainig Algorithm
$\mathcal{F}_{i}$ Fold $i$, used as validation set in the $i$-th iteration
$\mathcal{D}_{\text{train}}$ Training set formed by $K-1$ folds
@@
$\mathcal{M}_{i}$ Model trained in the $i$ iteration of cross-validation
@@
3:     $\mathcal{D}_{\text{train}}\leftarrow\bigcup_{j\neq i}\mathcal{F}_{j}$Combine $K-1$ folds for training}
@@ III-A Error Metrics
where the MCC value ranges from $-1$ to $+1$. 1 Represents perfect classification, 0 indicates random or uninformed classification, and $-1$ Denotes complete misclassification.
III-B Baseline Performance Without SMOTE-Tomek
The experimental results, detailed in Table IV and illustrated in Figure 3, reveal that the Linear SVM model consistently surpassed all other classifiers across the evaluated metrics. It achieved the highest performance, with a mean accuracy of 71.10%, precision of 68.25%, recall of 71.10%, F1-score of 66.47%, and an MCC of 0.5977.
Among the KNN models, configurations with k=3 and k=5 exhibited robust performance, with mean accuracies of 67.69% and 68.42%, respectively, and MCC values closely aligned with the Linear SVM, indicating their effectiveness on this dataset. The Gradient Boosting classifier also demonstrated competitive results, with a mean accuracy of 66.98% and an MCC of 0.5344, indicating its ability to model complex decision boundaries. Conversely, ensemble methods such as AdaBoost and Random Forest exhibited suboptimal performance, with AdaBoost showing particularly low precision of 28.21% and MCC of 0.1963, reflecting limitations in handling class imbalances. Traditional classifiers such as LR and DTs achieved moderate performance levels but were consistently outperformed by more advanced approaches like SVM and KNN.
TABLE IV: Performance For Each Model Without SMOTETomek
@@ III-B Baseline Performance Without SMOTE-Tomek
Figure 3: Cross validation performance of ML methods without SMOTE-Tomek.
III-C Enhanced Performance Using SMOTE-Tomek
Following the application of the SMOTE-Tomek preprocessing technique, the LR model demonstrated significant performance improvements, as shown in Figure 4 and Table V. The model achieved a mean accuracy of 76.16% ± 2.58%, a mean precision of 75.31% ± 4.30%, a mean recall of 76.16% ± 2.58%, a mean F1-score of 74.34% ± 2.93%, and a MCC of 0.67 ± 0.03. Performance was further enhanced, with accuracy improving by approximately 1%, through grid search optimization using C=10, L2 penalty, and saga solver. Here, C controls the trade-off between model complexity and training accuracy, with C=10 favoring a closer fit to the data. The L2 penalty regularizes model coefficients to prevent overfitting, while the saga solver is an efficient optimization algorithm designed for large datasets [32].
In comparison, Naive Bayes and Gradient Boosting also performed robustly, with mean accuracies exceeding 70% and MCC values of 0.62 and 0.59, respectively. Conversely, AdaBoost showed significant limitations, achieving a mean accuracy of just 16.10% and an MCC of 0.14, indicating its unsuitability for the current classification task.
Figure 4: Cross validation performance of machine learning method with SMOTE-Tomek.
TABLE V: Performance For Each Model With SMOTETomek
@@ III-C Enhanced Performance Using SMOTE-Tomek
Decision Tree 41.79 ± 12.44 63.89 ± 7.90 41.79 ± 12.44 40.43 ± 9.53 0.3363 ± 0.0780
Random Forest 59.03 ± 4.50 60.82 ± 6.96 59.03 ± 4.50 58.05 ± 5.39 0.4562 ± 0.0686
SVM Linear 74.20 ± 2.83 73.46 ± 4.58 74.20 ± 2.83 71.44 ± 3.26 0.6428 ± 0.0424
SVM RBF 60.27 ± 2.12 61.34 ± 5.40 60.27 ± 2.12 52.68 ± 3.09 0.4358 ± 0.0346
Naive Bayes 70.48 ± 2.28 75.04 ± 2.42 70.48 ± 2.28 71.17 ± 2.06 0.6286 ± 0.0280
Logistic Regression 76.16 ± 2.58 75.31 ± 4.30 76.16 ± 2.58 74.34 ± 2.93 0.6736 ± 0.0370
KNN (3 Neighbors) 57.79 ± 3.65 69.21 ± 3.19 57.79 ± 3.65 59.59 ± 3.78 0.5063 ± 0.0413
KNN (5 Neighbors) 51.60 ± 4.31 68.56 ± 3.76 51.60 ± 4.31 53.00 ± 5.20 0.4596 ± 0.0451
KNN (7 Neighbors) 46.34 ± 4.45 66.64 ± 4.05 46.34 ± 4.45 47.11 ± 5.69 0.4148 ± 0.0463
Gradient Boosting 70.90 ± 2.77 69.08 ± 4.15 70.90 ± 2.77 67.76 ± 3.56 0.5938 ± 0.0419
AdaBoost 16.10 ± 4.42 12.67 ± 6.97 16.10 ± 4.42 9.06 ± 3.07 0.1422 ± 0.0526
CatBoost 68.21 ± 3.32 68.34 ± 4.55 68.21 ± 3.32 66.37 ± 3.83 0.5633 ± 0.0492
III-D Impact of SMOTE-Tomek on Logistic Regression Interpretability
Understanding feature importance is critical in requirements classification tasks, as it provides actionable insights into the linguistic patterns that define requirement categories. Logistic regression (LR), as an interpretable model, offers a transparent mechanism for analyzing these relationships through its coefficients.
The coefficients presented in Table VI (LR with SMOTE-Tomek) and Table VII (LR without SMOTE-Tomek) highlight the influence of key features (words) in predicting requirement types. These coefficients quantify the strength and direction of association between words and their respective types. Positive coefficients indicate strong alignment with a particular requirement type, while the magnitude reflects the importance of the word in driving classification decisions. As an iterpretable model, LR benefits considerably from preprocessing techniques that address noise and class imbalance. SMOTE-Tomek enhances model performance, stabilizes, and refines coefficients.
The results reveal critical insights into the predictive influence of words and the effect of preprocessing on model behavior. For example, while the word ”seconds” in the ”Functional” type exhibited an extreme negative coefficient (-16.59) without preprocessing, SMOTE-Tomek balanced the dataset and reduced this coefficient to 6.55, ensuring a more consistent representation. Similarly, new influential words such as ”interface” in the ”Look-and-Feel” type emerged after preprocessing, indicating that SMOTE-Tomek enhances the visibility of class-distinguishing features.
The coefficients illustrate the inherent relationship between words and requirement types. High coefficients for words such as ”24x7” in ”Availability” and ”encrypted” in ”Security” highlight their criticality in these types. These insights can guide the development of keyword-based heuristics and inform domain-specific feature engineering.
Furthermore, Figure 5 illustrates the stabilizing effect of SMOTE-Tomek on absolute coefficient values, reducing variability and noise in the model.
TABLE VI: Top Words and Their Logistic Regression Coefficients
with SMOTE-Tomek
Class Top Words Coefficients
Functional allow, meeting, capability 6.55, 3.91, 3.83
Availability available, time, connected 8.67, 5.23, 4.91
Legal comply, regulations, case 8.32, 4.66, 3.88
Look-and-Feel interface, simulate, colors 5.46, 5.05, 4.37
Maintainability maintenance, new, releases 6.43, 5.41, 4.31
Operability server, database, microsoft 6.18, 5.08, 5.06
Performance seconds, let, performance 10.29, 7.03, 5.34
Scalability years, simultaneous, 000 5.88, 5.85, 5.14
Security access, data, security 7.05, 5.68, 5.19
Usability 90, training, easy 5.88, 5.06, 4.43
Fault Tolerance fault, goes, continue 7.20, 4.34, 3.96
Portability windows, portable, 2003 6.85, 5.97, 5.32
TABLE VII: Top Words and Their Logistic Regression Coefficients
without SMOTE-Tomek
Requirement Top Words Coefficients
Functional capability, allow, seconds 17.81, 17.69, -16.59
Availability 24x7, availability, connected 19.22, 14.67, 12.34
Legal comply, meet, regulations 20.25, 18.39, 15.36
Look-and-Feel simulate, colors, drill 18.84, 16.10, 13.74
Maintainability navigator, budget, season 14.93, 14.17, 10.90
Operability encoded, provides, distribution 14.13, 13.98, 13.36
Performance seconds, let, performance 28.51, 24.79, 18.85
Scalability simultaneous, 000, capable 13.29, 12.56, 11.82
Security encrypted, access, role 17.40, 17.03, 15.04
Usability training, 90, help 15.65, 14.39, 14.38
Fault Tolerance loss, fault, downloaded 21.61, 19.65, 16.55
Portability portable, mobile, major 15.97, 12.39, 12.02
Figure 5: Comparison of top word coefficients with and without SMOTE-Tomek (absolute values).
III-E Discussion

### v2

@@ Abstract
Requirements Engineering , Class Imbalance , SMOTE-Tomek , Imbalanced Datasets , Requirements Classification , Natural Language Processing , Synthetic Oversampling , Text-Based Classification , Supervised Learning
1 Introduction
@@
Requirements engineering (RE) is a cornerstone of the software development lifecycle, translating stakeholder needs into actionable system specifications [kotonya1998requirements, lanfear2024enhancing, cai2024unified]. Accurate and efficient requirement classification is essential to ensure project clarity, prioritization, and goal alignment. Requirements are typically categorized into functional, non-functional, and subtypes [dias2020software, quba2021software, winkler2016automatic, binkhonain2023machine, wei2023data]. A survey revealed that over 60% of failed projects neglected non-functional requirements, underscoring the critical importance of effective classification [bajpai2012non]. Despite its significance, requirement classification remains a largely manual process, prone to inconsistencies, inefficiencies, and scalability challenges [van2000handling].
@@
Over the years, researchers have explored various approaches to automate requirement classification. Early rule-based systems were among the first attempts, offering interpretable but inflexible solutions that were labor-intensive to maintain [spanoudakis2004rule]. The advent of machine learning (ML) brought more adaptive models, demonstrating moderate success, particularly in structured datasets [jordan2015machine, dias2020software, binkhonain2023machine]. However, these methods often face challenges with class imbalance, a prevalent issue in real-world datasets, including the widely used PROMISE dataset for requirements engineering [lima2019software]. This imbalance skews model predictions toward majority classes, undermining performance on minority classes.
@@
Over the past decade, deep learning (DL) has achieved transformative breakthroughs across multiple domains, including natural language processing (NLP), computer vision, and time series analysis. These advancements are primarily driven by the unprecedented representational power of deep neural networks, which leverage millions, and in some cases billions, of trainable parameters to extract and model intricate, high-dimensional patterns from large-scale datasets [lecun2015deep].
@@
Equally significant are the advances in time series analysis, where DL models have demonstrated remarkable capabilities in tasks such as motion sensing classification and physical quantity estimation [freydin2022learning, freydin2024deep, or2024carspeednet, or2024transformer]. These innovations highlight DL’s ability to capture temporal dependencies and complex properties in sequential data.
@@
In computer vision, architectures such as ResNet [he2016deep] and Vision Transformers (ViT) [dosovitskiy2020image] have revolutionized image classification, object detection, and semantic segmentation, achieving state-of-the-art accuracy while enabling applications that were previously unattainable.
@@
In NLP, DL models such as BERT [devlin2018bert] and GPT-3 [brown2020language] have set new standards in tasks like language translation, text summarization, and question answering, surpassing traditional methods and enabling more nuanced understanding and generation of human language [vaswani2017attention, luo2022prcbert, subahi2023bert, belani2019requirements, navarro2017towards]. These models excel in NLP tasks due to their ability to capture contextual relationships within text using self-attention mechanisms. They are particularly effective at modeling semantic intricacies in textual data, enabling strong performance in complex classification tasks. However, these methods require not only substantial computational resources for fine-tuning and inference but also access to massive amounts of high-quality data to achieve their full potential. This reliance on extensive datasets and powerful hardware often makes them less practical for small-scale or resource-constrained projects, where data availability and computational capacity may be limited.
@@
To address these challenges, this study integrates the Synthetic Minority Oversampling Technique (SMOTE) [chawla2002smote] with Tomek Links (SMOTE-Tomek) [batista2004study], providing a robust preprocessing solution for imbalanced datasets. SMOTE oversamples minority classes by generating synthetic samples through interpolation, effectively enhancing the diversity of underrepresented classes while mitigating overfitting, a common issue with random oversampling methods. Complementing this, Tomek Links identifies and removes noisy or borderline samples, improving class separability by reducing overlapping data points between classes. This dual-action preprocessing strategy results in cleaner and more balanced training datasets, enabling ML models to better learn decision boundaries for minority classes.
@@
In RE classification, SMOTE-Tomek addresses both class imbalance and noise, making it particularly suitable for the PROMISE dataset, where minority classes are significantly underrepresented. While previous studies have demonstrated the efficacy of SMOTE-Tomek for addressing class imbalance in the PROMISE dataset [abbas2024enscl], this research extends its application by integrating it into a stratified K-fold cross-validation framework. This ensures the validation folds remain untouched by resampling, preserving the original class distribution and enabling rigorous evaluation of the model’s generalization capabilities. Furthermore, this study explores several classical ML models, evaluates their performance on the PROMISE dataset, and highlights the potential of lightweight solutions for efficient requirement classification.
@@ 2 Learning Method
2.1 Dataset
The study utilizes the expanded PROMISE dataset, a diverse collection of 969 software requirements extracted from Software Requirements Specification (SRS) documents using the Google search engine [lima2019software]. The dataset contains 444 functional requirements (45.8%) and 525 non-functional requirements (54.2%), distributed across 12 distinct categories. These categories represent a spectrum of requirements in software engineering, such as Security (SE), Usability (US), Portability (PO), and Performance (PE). Despite its comprehensiveness, the dataset exhibits notable class imbalance. For example, the Portability category is severely underrepresented, comprising only 12 requirements (1.24%), while the most prevalent non-functional class, Security, includes 125 requirements (12.9%). This disparity is further amplified in the broader classification between functional and non-functional requirements, which affects the performance and generalizability of ML models. Figure 1 illustrates the distribution of the dataset across all requirement types, highlighting the inherent imbalance.
@@ 2.1 Dataset
Table 1: Example requirements and their corresponding types from the PROMISE dataset
@@
Users should be able to access their streaming movies in under two clicks after logging into the website. Usability (US)
@@
2.2 Pre-Processing
Transforming unstructured text into a numerical representation suitable for ML algorithms is a fundamental preprocessing step in many NLP tasks. The PROMISE dataset comprises individual sentences, rather than full documents. The common Term Frequency-Inverse Document Frequency (TF-IDF) remains a highly effective vectorization technique [ramos2003using, robertson2004understanding]. TF-IDF quantifies the importance of a term within a sentence while considering its prevalence across the entire dataset of sentences. By emphasizing terms that are frequent within a specific sentence but rare across the dataset, TF-IDF captures features that are contextually meaningful and relevant to the classification task. Despite the brevity of the textual units, TF-IDF preserves the semantic significance of terms.
2.3 Class Imbalance Challenge
Class imbalance poses a critical challenge in training ML models, as they tend to prioritize majority classes, resulting in poor generalization and suboptimal performance on minority classes [johnson2019survey]. Such biases can significantly impact the usability and reliability of classification models in practical applications, where correctly identifying minority classes is often essential. Balancing the training set ensures that the model learns to give equal attention to all classes, preventing it from being overly biased toward majority classes.
2.4 SMOTE-Tomek
@@
In practical applications of NLP tasks, where textual data is complex and often imbalanced, SMOTE-Tomek emerges as a robust and effective solution. The synthetic examples generated by SMOTE are not direct textual entities but instead constitute feature vectors that represent the underlying characteristics of the minority class, as presented in Table 2. These vectors serve as abstract representations, enabling the model to internalize the structural properties of underrepresented classes without requiring additional annotated textual data.
@@
Table 2: Examples of resampled requirement representations using SMOTE-Tomek
Requirement Representation Requirement Type
based, established, lead, operate, physical, process, product, run, server, service, shall, structure, washing, web Operability (O)
changes, edit, information, just, modify, normal, personal, read, securely, server, transmission, transmitted, users Security (SE)
entry, shown Functional (F)
initiator, mediator, messages, request, response, shall, user Functional (F)
accommodate, compensatory, data, failures, fault, items, large, product, recovery, robust, routing, shall, technique, tolerance, transaction, using Fault Tolerance (FT)
explanatory, intuitive, self, shall Usability (US)
control, details, employees, site, view Functional (F)
browsers, ce, compatible, environments, expected, interface, major, operating, order, palm, product, run, standards, systems, usable, variety, web, wide, windows Portability (PO)
800x600, 95, appropriately, correct, corrected, display, feel, higher, implementation, incorrect, intranet, look, notification, pages, prior, provide, remaining, resolutions, shall, uniform, web, week Look-and-Feel (LF)
able, alongside, days, environment, function, installed, java, operating, product, runtime, server, shall, software Operability (O)
2.5 Stratified K-fold Cross-Validation Method
@@
2.6 Trainig Algorithm
The proposed method, as outlined in Algorithm 1, emphasizes the integration of stratified K-fold cross-validation to ensure class proportions are preserved across folds. Importantly, SMOTE-Tomek is applied solely to the training folds to prevent data leakage, thereby preserving the integrity of validation metrics. The algorithm also includes final training on the entire balanced dataset to fully utilize the available data. Table 3 presents the notations and their corresponding definitions as utilized in the algorithm.
Table 3: Notations and definitions
@@ 2.6 Trainig Algorithm
$\mathcal{F}_{i}$ Fold $i$ used as the validation set in the $i$-th iteration
$\mathcal{D}_{\text{train}}$ Training set formed by K1K-1 folds
@@
$\mathcal{M}_{i}$ Model trained in the $i$-th cross-validation iteration
@@
3:  $\mathcal{D}_{\text{train}}\leftarrow\bigcup_{j\neq i}\mathcal{F}_{j}$ {Combine K1K-1 folds for training}
@@ 3.1 Error Metrics
where the MCC value ranges from 1-1 to +1+1. 1 Represents perfect classification, 0 indicates random or uninformed classification, and 1-1 Denotes complete misclassification.
3.2 Baseline Performance Without SMOTE-Tomek
The baseline experimental results are summarized in Table 4 and visualized in Figure 3. Among all evaluated classifiers, the linear SVM achieved the strongest overall performance, attaining a mean accuracy of 71.10%, precision of 68.25%, recall of 71.10%, F1-score of 66.47%, and an MCC of 0.5977. This indicates a favorable balance between predictive accuracy and robustness under class imbalance.
KNN classifiers with $k=3$ and $k=5$ also demonstrated competitive performance, with mean accuracies of 67.69% and 68.42%, respectively, and MCC values close to that of the linear SVM. These results suggest that local neighborhood–based methods can capture discriminative structure in the PROMISE dataset despite class imbalance. Gradient Boosting further exhibited solid performance, achieving a mean accuracy of 66.98% and an MCC of 0.5344, reflecting its ability to model non-linear decision boundaries.
In contrast, ensemble methods such as Random Forest and AdaBoost showed substantially weaker results. AdaBoost, in particular, suffered from severe degradation, with a mean precision of only 28.21% and an MCC of 0.1963, highlighting its sensitivity to imbalanced class distributions. Classical models such as Logistic Regression and Decision Trees achieved moderate performance but were consistently outperformed by SVM and KNN variants in the absence of imbalance-aware preprocessing.
Table 4: Performance for each model without SMOTE-Tomek
@@ 3.2 Baseline Performance Without SMOTE-Tomek
Figure 3: Cross-validation performance of machine learning models without SMOTE-Tomek preprocessing.
3.3 Enhanced Performance Using SMOTE-Tomek
The impact of SMOTE-Tomek preprocessing on classification performance is presented in Table 5 and Figure 4. Following the application of SMOTE-Tomek, Logistic Regression exhibited the most pronounced improvement across all evaluation metrics, achieving a mean accuracy of 76.16% $\pm$ 2.58%, precision of 75.31% $\pm$ 4.30%, recall of 76.16% $\pm$ 2.58%, F1-score of 74.34% $\pm$ 2.93%, and an MCC of 0.6736 $\pm$ 0.0370.
This represents a substantial gain relative to the baseline Logistic Regression results reported in Table 4. Performance was further improved by approximately one percentage point through hyperparameter optimization using $C=10$, an L2 regularization penalty, and the SAGA solver, which is well suited for large-scale linear models [defazio2014saga].
Several additional classifiers benefited from SMOTE-Tomek preprocessing. Naive Bayes and Gradient Boosting achieved mean accuracies exceeding 70%, with MCC values of 0.6286 and 0.5938, respectively, indicating improved robustness to class imbalance. In contrast, AdaBoost remained highly sensitive to the resampled data distribution, achieving a mean accuracy of only 16.10% and an MCC of 0.1422, confirming its unsuitability for this task under severe imbalance conditions.
Figure 4: Cross-validation performance of machine learning models with SMOTE-Tomek preprocessing.
Table 5: Performance for each model with SMOTE-Tomek
@@ 3.3 Enhanced Performance Using SMOTE-Tomek
Decision Tree 41.79 $\pm$ 12.44 63.89 $\pm$ 7.90 41.79 $\pm$ 12.44 40.43 $\pm$ 9.53 0.3363 $\pm$ 0.0780
Random Forest 59.03 $\pm$ 4.50 60.82 $\pm$ 6.96 59.03 $\pm$ 4.50 58.05 $\pm$ 5.39 0.4562 $\pm$ 0.0686
SVM (Linear) 74.20 $\pm$ 2.83 73.46 $\pm$ 4.58 74.20 $\pm$ 2.83 71.44 $\pm$ 3.26 0.6428 $\pm$ 0.0424
SVM (RBF) 60.27 $\pm$ 2.12 61.34 $\pm$ 5.40 60.27 $\pm$ 2.12 52.68 $\pm$ 3.09 0.4358 $\pm$ 0.0346
Naive Bayes 70.48 $\pm$ 2.28 75.04 $\pm$ 2.42 70.48 $\pm$ 2.28 71.17 $\pm$ 2.06 0.6286 $\pm$ 0.0280
Logistic Regression 76.16 $\pm$ 2.58 75.31 $\pm$ 4.30 76.16 $\pm$ 2.58 74.34 $\pm$ 2.93 0.6736 $\pm$ 0.0370
KNN (k=3) 57.79 $\pm$ 3.65 69.21 $\pm$ 3.19 57.79 $\pm$ 3.65 59.59 $\pm$ 3.78 0.5063 $\pm$ 0.0413
KNN (k=5) 51.60 $\pm$ 4.31 68.56 $\pm$ 3.76 51.60 $\pm$ 4.31 53.00 $\pm$ 5.20 0.4596 $\pm$ 0.0451
KNN (k=7) 46.34 $\pm$ 4.45 66.64 $\pm$ 4.05 46.34 $\pm$ 4.45 47.11 $\pm$ 5.69 0.4148 $\pm$ 0.0463
Gradient Boosting 70.90 $\pm$ 2.77 69.08 $\pm$ 4.15 70.90 $\pm$ 2.77 67.76 $\pm$ 3.56 0.5938 $\pm$ 0.0419
AdaBoost 16.10 $\pm$ 4.42 12.67 $\pm$ 6.97 16.10 $\pm$ 4.42 9.06 $\pm$ 3.07 0.1422 $\pm$ 0.0526
CatBoost 68.21 $\pm$ 3.32 68.34 $\pm$ 4.55 68.21 $\pm$ 3.32 66.37 $\pm$ 3.83 0.5633 $\pm$ 0.0492
3.4 Discussion